OpenStudy (anonymous):

Evaluate each expression for the given value of the variable 14-b for b= -3

OpenStudy (anonymous):

14-(-3) 14+3=17

OpenStudy (anonymous):

Thank

OpenStudy (anonymous):

np

OpenStudy (anonymous):

-7 -q for q =-15

OpenStudy (anonymous):

-7+15=

OpenStudy (anonymous):

-5-12=

OpenStudy (anonymous):

A submarine cruising at 27 m below sea level or -27 m descends 14 m

OpenStudy (dangerousjesse):

-|27+14|=?

OpenStudy (anonymous):

Then what do I do

OpenStudy (dangerousjesse):

Just solve the equation to get the answer, you do nothing else with it.

OpenStudy (anonymous):

😏

OpenStudy (dangerousjesse):

A random square?

OpenStudy (dangerousjesse):

No, I don't think that's the answer...